I don't get this one, I get stuck on the 3rd step.
(x+9)^2=18
=x^2+81-18=0
=x^2+63=0
I don't know what to do after this since there is no square root of 63 that is a whole number. help?

Answer by ankor@dixie-net.com(22740)

(x+9)^2 = 18
If you FOIL(x+9)(x+9),you should get
x^2 + 18x + 81 = 18
x^2 + 18x + 63 = 0
this will not factor, use the quadratic formula to find x